In an electric shaver, the blade moves back and forth over a distance of 2.0 mm. The motion is simple harmonic, with frequency 110 Hz.

(a) Find the amplitude.
mm
(b) Find the maximum blade speed.  
m/s
(c) Find the maximum blade acceleration.
m/s2
Answer #1

total distance = 2A = 2 mm

Amplitude = A = 2/ 2 = 1 mm = 0.001 m

b) w = 2 pie * f = 2 *3.14 * 110 = 690.8 rad/sec

maximum speed = Vmax = A W = 0.001 * 690.8 = 0.69 m/sec

c) amax = w^2 * A = 690.8^2 * 0.001= 477 m/sec^2